OpenStudy (ellamoyseyuk):

simplify. write answers with negative exponents. (abc)^-3 c^2 b ------------ a^-4 bc^2 a

OpenStudy (ellamoyseyuk):

i dont know how to turn this into a negative exponent. can you help me @agent0smith

OpenStudy (agent0smith):

Simplify it as normal first, then worry about the negative exponents.

OpenStudy (ellamoyseyuk):

i already did it where it say up top that this is the step before last

OpenStudy (agent0smith):

Completely simplify it. That's not even close.

OpenStudy (ellamoyseyuk):

wait that was the wrong step sorry

OpenStudy (agent0smith):

Completely simplify it, then move terms to the numerator (or denominator) to make them negative. If you know how to make a negative exponent into a positive exponent, then you can do the reverse.

OpenStudy (ellamoyseyuk):

\frac{ a ^{-3}b ^{2}c ^{-1}} }{a ^{-3}bc ^{2} }

OpenStudy (agent0smith):

I can tell that's still not completely simplified...

OpenStudy (ellamoyseyuk):

ac^3/b^-3

OpenStudy (ellamoyseyuk):

is that right or no?

OpenStudy (ellamoyseyuk):

but it is supposed to be negative and were did the a go?

OpenStudy (ellamoyseyuk):

(a) go

OpenStudy (agent0smith):

The \(a\) canceled out. It's easy to now make them negative. If you know how to make a negative exponent into a positive exponent, then you can do the reverse of that.

OpenStudy (ellamoyseyuk):

1 / b^-3 c^-3

OpenStudy (agent0smith):

\[\Large x^{-b} = \frac{ 1 }{ x^b }\]shouldn't be hard to do that in reverse... start from the right side, go back to the left

OpenStudy (ellamoyseyuk):

b^-3 = 1/b^3 c^-3=1/c^3

OpenStudy (niksbhalla14mar):

The question is (abc)^-3 c^2 b ------------ a^-4 bc^2 a Lets bring the division as multiplication , so we would get the signs as opposite. So it becomes (abc)^-3 . c^2 . b . a^4 . b^-1 . c^-2 . a^-1 Further solving we get a^(-3+4-1) . b^(-3+1-1) . c^(-3+2-2) = a^0 . b^-3 . c^-3 =1/(b^3.c^3) is the final answer. simply solve for all the powers of a , b & c individually.